# Leveraging Distribution Matching to Make Approximate Machine Unlearning Faster

### v2

@@ Abstract
Approximate machine unlearning (AMU) enables models to ‘forget’ specific training data through specialized fine-tuning on a retained dataset subset. However, processing this retained subset still dominates computational runtime, while reductions of epochs also remain a challenge. We propose two complementary methods to accelerate classification-oriented AMU. First, Blend, a novel distribution-matching dataset condensation (DC), merges visually similar images with shared blend-weights to significantly reduce the retained set size. It operates with minimal pre-processing overhead and is orders of magnitude faster than state-of-the-art DC methods. Second, our loss-centric method, Accelerated-AMU (A-AMU), augments the unlearning objective to quicken convergence. A-AMU achieves this by combining a steepened primary loss to expedite forgetting with a differentiable regularizer that matches the loss distributions of forgotten and in-distribution unseen data. Our extensive experiments demonstrate that this dual approach of data and loss-centric optimization dramatically reduces end-to-end unlearning latency across both single and multi-round scenarios, all while preserving model utility and privacy. To our knowledge, this is the first work to systematically tackle unlearning efficiency by jointly designing a specialized dataset condensation technique with a dedicated accelerated loss function. Code is available at https://github.com/algebraicdianuj/DC_Unlearning.
@@ 1 Introduction
Machine unlearning (MU) seeks to expunge specific training points from a deployed model. Exact MU [2, 1] offers perfect deletion guarantees by fully retraining, often via data‐sharding schemes, but remains prohibitively slow for modern networks. Approximate MU (AMU) [13, 15, 6, 9, 25, 11] replaces such guarantees with speed, typically minimizing in iterative or one-step settings:
$\alpha\,\underbrace{\textbf{Main Loss}}_{\text{on retained data}}\,+\,\beta\,\underbrace{\textbf{Forgetting Reg.}}_{\text{drives deletion}}$
@@
We address this by realizing reduction of the retained dataset via specialized speed-focused dataset condensation [30]. Building upon distribution-matching dataset condensation [29], we propose Blend [^1] This method is a revision of our previous work available at https://arxiv.org/abs/2402.00195., that learns a compact blend-weight vector to synthesize a tiny set of synthetic datasets that preserve training utility. Integrating this into AMU pipeline shrinks the retained set and accelerates unlearning without sacrificing accuracy or privacy, with a minor overhead. We further push for loss-centric optimization through Accelerated-AMU (A-AMU) loss that steepens the main loss landscape, alongside loss distribution matching formulation (as a MIA minimizer), enabling rapid convergence to the unlearned model. Our approach demonstrates substantial speedups and very strong unlearning compared to baseline SOTA unlearning methods. We summarize our contributions as follows:
@@ 3.2 Blend Dataset Condensation
Under Algorithm 1, we divide images of each class of D into $k$ sub-classes using $k$-means clustering over the feature representations, extracted by a randomly initialized lightweight feature extractor $\psi_{\vartheta}$ (much smaller than the pretrained model to be unlearned). The resulting partition of the dataset can be written as:
[…]
@@
Algorithm 1 $k$-means Partitioning ($\mathbf{D}=\{I_{p}\}_{p}$)
@@
2:$\mathcal{C}\leftarrow\varnothing$triangleright Initialize image clusters
@@
4:    $\{I_{a}\}_{a=1}^{N_{c}}\leftarrow\{I_{p}\mid y_{p}=c\}$triangleright Collect class images
5:    $\{\psi_{\vartheta}(I_{a})\}_{a=1}^{N_{c}}$triangleright Compute feature vectors
6:    $\{\psi_{\vartheta}(I_{a})\}_{a=1}^{N_{c}}\xrightarrow{k\text{-means}}\bigl{\{}\{\psi_{\vartheta}(I_{b\kappa})\}_{b=1}^{n_{\kappa}}\bigr{\}}_{\kappa=1}^{k}$triangleright Run $k$-means
7:    $\bigl{\{}\{\psi_{\vartheta}(I_{b\kappa})\}_{b=1}^{n_{\kappa}}\bigr{\}}_{\kappa=1}^{k}\to\bigl{\{}\{I_{b\kappa}\}_{b=1}^{n_{\kappa}}\bigr{\}}_{\kappa=1}^{k}$triangleright Map to images
8:    $\mathcal{C}\leftarrow\mathcal{C}\cup\bigl{\{}\{I_{b\kappa}\}_{b=1}^{n_{\kappa}}\bigr{\}}_{\kappa=1}^{k}$triangleright Add to global set
@@
10:$\mathcal{C}=\bigl{\{}\{\{I_{b\kappa c}\}_{b=1}^{n_{\kappa c}}\}_{\kappa=1}^{k}\bigr{\}}_{c=1}^{C}$triangleright Full cluster set
@@
12:return $\{\{I_{ij}\}_{i=1}^{n_{j}}\}_{j=1}^{k\times C}$triangleright Partitioned D dataset
@@
[…]
@@ 3.3 Creating the Reduced Retain Dataset
Algorithm 2 systematically separates the retain images (using partitioned training set and $\mathbf{F}$) into:
@@
[…]
@@
2:$\phi(j)=\begin{cases}-1,&\mathbf{F}\cap\{I_{ij}\}_{j=1}^{k\times C}=\emptyset\\ j,&\text{otherwise}\end{cases}$triangleright indexing not $\mathbf{F}$ clusters
3:$\Phi(i)=\begin{cases}i,&\mathbf{F}\cap\{I_{ij}\}_{i=1}^{n_{j}}\\ -1,&\text{otherwise}\end{cases}$triangleright indexing within $\mathbf{F}$ clusters
4:$\mathbf{R}_{\text{free}}=\{\{I_{ij}\}_{i=1}^{n_{j}}\}_{j=\phi(1)}^{\phi(kC)}$triangleright $\mathbf{R}$ subset dissimilar to $\mathbf{F}$
5:$\mathbf{R}_{\text{residual}}=\{\{I_{ij}\}_{i=\Phi_{j}(1)}^{\Phi_{j}(n_{j})}\}_{j=1}^{kC}$triangleright $\mathbf{R}$ subset similar to $\mathbf{F}$
@@ 3.4 Accelerated Approximate Unlearning
Carlini et al. [3] report that per-sample classification losses of deep networks are strongly non-Gaussian. They considered to “Gaussianise” losses by ad-hoc monotone maps (e.g. combination of $\exp$ and $\log$ maps), without guaranteeing that the output is normally distributed. We instead apply the probability–integral transform (PIT): for any continuous variable $X$ with CDF $F_{X}$, $Z=\Phi^{-1}\!\bigl{(}F_{X}(X)\bigr{)}$ is standard normal. Substituting $F_{X}$ by a differentiable empirical surrogate yields an end-to-end differentiable pipeline whose transformed losses converge to $\mathcal{N}(0,1)$. Following this motivation, we intend to transform arbitrary random samples, which may not be following Gaussian distribution towards samples following standard normal distribution. Let $x=(x_{1},\ldots,x_{N})\;\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\;F$, such that $F$ is an unknown continuous CDF. The empirical CDF at $x_{i}$ is $\hat{F}_{N}(x_{i})=\tfrac{1}{N}\sum_{j=1}^{N}\mathbf{1}\{x_{j}\leq x_{i}\}.$ Replacing the indicator by the temperature–controlled sigmoid $\sigma_{K}(z)=(1+e^{-Kz})^{-1},\;K>0,$ gives the differentiable estimator
@@
Because each entry of $\mathbf{Q}^{(K)}(x)$ lies in $(0,1)$, apply the probit (quantile) function $\Phi^{-1}$ of the standard normal distribution $\mathcal{N}(0,1)$ (mean 00, variance $1$) to obtain [22]
$\mathbf{z}^{(K)}=\Phi^{-1}\!\bigl{(}\mathbf{Q}^{(K)}(x)\bigr{)}\in\mathbb{R}^{N}.$ (5)
Thus any sample $x$ from a continuous distribution is converted to a sample whose coordinates converge to i.i.d. standard normals as $K\to\infty$: $\mathbf{z}^{(K)}\;\xrightarrow{K\to\infty}\;\Phi^{-1}\!\bigl{(}\hat{F}_{N}(x)\bigr{)}.$ In practice one chooses $K$ finite but large to balance smooth gradients (for optimization) with closeness to the empirical CDF.
@@ 3.4.1 Transforming losses of model to standard normal
Applying the vector transformation in (5) to the loss vectors yields $\mathbf{Z}_{F}^{(K)}=\Phi^{-1}\!\bigl{(}\mathbf{Q}^{(K)}(\bm{\ell}_{F})\bigr{)},\qquad\mathbf{Z}_{T}^{(K)}=\Phi^{-1}\!\bigl{(}\mathbf{Q}^{(K)}(\bm{\ell}_{T})\bigr{)},$ where $\mathbf{Q}^{(K)}$ is the differentiable CDF estimator from (4) and $\Phi^{-1}$ is the probit of the standard normal distribution $\mathcal{N}(0,1)$. As $K\to\infty$ the mapping converges to the empirical probit, and this allows use to play with transformed losses in distribution matching manner. As emphasize previously, practice we keep $K$ sufficiently large to preserve differentiability.
@@ 3.4.2 Membership-Inference Attack (MIA) Loss
$\mathcal{L}_{\mathrm{MMD}}(\theta)$ $=\mathbb{E}_{z,z^{\prime}\sim\mathbf{Z}_{F}^{(K)}}\!\bigl{[}\curlywedge(z,z^{\prime})\bigr{]}+\mathbb{E}_{\tilde{z},\tilde{z}^{\prime}\sim\mathbf{Z}_{T}^{(K)}}\!\bigl{[}\curlywedge(\tilde{z},\tilde{z}^{\prime})\bigr{]}$ (6)
 $\quad-\,2\,\mathbb{E}_{z\sim\mathbf{Z}_{F}^{(K)},\,\tilde{z}\sim\mathbf{Z}_{T}^{(K}}\!\bigl{[}\curlywedge(z,\tilde{z})\bigr{]}.$
@@ 3.4.3 Accelerated (A-) AMU Objective
Replacing $\mathcal{L}$ by its square $\mathcal{L}^{2}$ multiplies the gradient magnitude by $2\,\mathcal{L}$. Consequently, gradient descent updates make larger progress early in training. Large errors on $\mathbf{R}$ are penalized disproportionately, while small errors become less important in course of optimization. So the model departs more rapidly from the decision boundaries, which it learned earlier during training schedule. In order to prevent overfitting on singular retain dataset samples for fixed learning rate, we take square of average loss per batch, instead of per sample squaring. This induces stability on choice of learning rate. Infact, in all of our experiments, we set a fixed learning rate for our approach on all models and datasets (including other hyperparameters), which is in strong contrast from other SOTA AMU methods that require hectic finetuning on each setting.
@@
Combining the controlled instability from steeper main loss, around decision boundaries, with reduction of training trajectories towards absolute unlearning solutions, forms a strong acceleration of unlearning. As would seen later in our experiments, A-AMU objective minimization achieves competitive MIA suppression and retain accuracy after $\sim 1\,\text{epoch}$.
@@ 4.1 Experimental Setup
We evaluate our methods on CIFAR-10, SVHN, and CINIC-10 datasets using ResNet-18 and ResNet-50 models. Our approach is benchmarked against six methods: Retraining, Catastrophic Forgetting (CF), Bad Distillation [6], $\mathbf{L_{1}}$-Sparsity [13], Pruning [13], and SCRUB [15]. For a fair comparison, implementations were adapted from their official repositories, and hyperparameters were rigorously fine-tuned for each model and dataset. [^2] Complete hyperparameter configurations are available in the ’hyperparameters’ folder of our code repository: https://github.com/algebraicdianuj/DC_Unlearning. We categorize these baselines into three tiers for analysis: naive (Retraining, CF), non-naive ($\mathbf{L_{1}}$-sparse, Pruning, SCRUB, Bad Distillation), and our proposed accelerated (A-) methods.
@@
Random Class Forgetting (Condensation Time for 1000 IPC: $47.7812\pm 1.3323$ s)
Retraining $48.61\pm 1.85$ $49.49\pm 2.38$ $100.00\pm 0.00$ $68.45\pm 2.51$ $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $79.96\pm 1.10$ $56.46\pm 2.81$ $3580.71\pm 94.46$ $705.82\pm 0.30$
CF $51.07\pm 3.71$ $57.87\pm 2.74$ $100.00\pm 0.00$ $97.89\pm 0.18$ $00.00\pm 0$ 0.00 $1.65\pm 1.73$ $81.28\pm 1.33$ $79.11\pm 1.89$ $1789.40\pm 47.15$ $352.38\pm 0.74$
Bad Distillation $57.97\pm 5.13$ $62.99\pm 6.81$ $97.31\pm 1.59$ $99.69\pm 0.09$ $00.00\pm 0$ 0.00 $0.01\pm 0.03$ $77.68\pm 2.36$ $80.64\pm 1.16$ $1370.90\pm 34.32$ $572.67\pm 0.69$
$L_{1}$ Sparse $49.18\pm 2.20$ $58.95\pm 4.72$ $98.52\pm 0.55$ $98.54\pm 0.32$ $00.00\pm 0$ 0.00 $4.03\pm 3.58$ $77.61\pm 2.12$ $79.61\pm 1.49$ $1446.43\pm 6.34$ $288.65\pm 0.31$
SCRUB $50.00\pm 0.00$ $50.00\pm 0.00$ $15.50\pm 3.00$ $11.14\pm 0.06$ $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $13.88\pm 2.72$ $10.05\pm 0.09$ $759.90\pm 9.19$ $198.01\pm 0.21$
Pruning $48.37\pm 2.43$ $50.56\pm 0.81$ $97.10\pm 1.14$ $99.08\pm 0.63$ $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $76.85\pm 1.84$ $69.80\pm 3.61$ $722.06\pm 5.91$ $143.73\pm 0.23$
A-CF $52.10\pm 1.50$ $50.99\pm 0.67$ $97.75\pm 1.50$ $95.90\pm 2.73$ $00.00\pm 0$ 0.00 $0.10\pm 0.19$ $77.03\pm 1.08$ $75.90\pm 3.82$ $218.30\pm 0.26$ $44.54\pm 1.80$
10% Uniform Forgetting (Condensation Time for 1000 IPC: $34.1780\pm 0.7403$ s)
@@
Unlearning performance is assessed across five key metrics: Test Accuracy ($\uparrow$), Retain Accuracy ($\uparrow$), Forget Accuracy ($\downarrow$), Unlearning Time ($\downarrow$), and a MIA Score ($\downarrow$) for privacy. The MIA score is derived from a powerful white-box LiRA attack [3], providing a robust measure of privacy risk. For our Blend method, the $k$-means partitioning feature extractor is a lightweight CNN, 35×35$\times$35 × and 74×74$\times$74 × smaller than ResNet-18 and ResNet-50 respectively, ensuring its preprocessing overhead remains minimal. All experiments were conducted on a single NVIDIA RTX 3070 GPU (8GB VRAM), except for experiments in Sections 4.4 and 4.5, which were performed on a laptop RTX 3070 (8GB VRAM) due to computational resource availability during the extended experimental phase.
@@
We categorize baselines into three tiers: naive (Retraining, CF), non-naive ($\mathbf{L_{1}}$-sparse, Pruning, SCRUB, Bad Distillation), and our accelerated (A-) which can contain acceleration of all naive and non-naive AMU methods.

### v3

@@ Abstract
Approximate machine unlearning (AMU) enables models to ‘forget’ specific training data through specialized fine-tuning on a retained (and forget) subset of training set. However, processing this large retained subset still dominates computational runtime, while reductions of unlearning epochs also remain a challenge. In this paper, we propose two complementary methods to accelerate arbitrary classification-oriented AMU method. First, Blend, a novel distribution-matching dataset condensation (DC), merges visually similar images with shared blend-weights to significantly reduce the retained set size. It operates with minimal pre-processing overhead and is orders of magnitude faster than state-of-the-art DC methods. Second, our loss-centric method, Accelerated-AMU (A-AMU), augments the AMU objective to quicken convergence. A-AMU achieves this by combining a steepened primary loss to expedite forgetting with a differentiable regularizer that matches the loss distributions of forgotten and in-distribution unseen data. Our extensive experiments demonstrate that this dual approach of data and loss-centric optimization dramatically reduces end-to-end unlearning latency across both single and multi-round scenarios, all while preserving model utility and privacy. To our knowledge, this is the first work to systematically tackle unlearning efficiency by jointly designing a specialized dataset condensation technique with a dedicated accelerated loss function. Code is available at https://github.com/algebraicdianuj/DC_Unlearning.
@@ 1 Introduction
Machine unlearning (MU) seeks to expunge specific training points from a deployed model. Exact MU [2, 1] offers perfect deletion guarantees by fully retraining, often via data‐sharding schemes, but remains prohibitively slow for modern networks. Approximate MU (AMU) [13, 15, 6, 9, 25, 11] replaces such guarantees with speed, by achieving following goal in one-step or iterative settings:
$\min_{\text{model parameters}}\alpha\,\underbrace{\textbf{Main Loss}}_{\text{on retained data}}\,+\,\beta\,\underbrace{\textbf{Forgetting Reg.}}_{\text{drives deletion}}$ (*)
@@
We address this by realizing reduction of the retained dataset via specialized speed-focused dataset condensation [30]. Building upon distribution-matching dataset condensation [29], we propose Blend [^1] This method is a revision of our previous work available at https://arxiv.org/abs/2402.00195., that learns a compact blend-weight vector to synthesize a tiny set of synthetic datasets that preserve training utility. Integrating this into AMU pipeline shrinks the retained set and accelerates unlearning without sacrificing accuracy or privacy, with a minor overhead. We further push for loss-centric optimization through Accelerated-AMU (A-AMU) loss that preserves the template (*1), except it steepens the main loss at per sample (or batch) level, and regularizes it with a distribution matching formulation of a membership inference attack (MIA) minimizer. This enables rapid convergence to the unlearned model. Our approach demonstrates substantial speedups and very strong unlearning compared to baseline SOTA AMU methods. We summarize our contributions as follows:
@@ 3.2 Blend Dataset Condensation
Using Algorithm 1, we divide images of each class of D into $k$ sub-classes using $k$-means clustering over the feature representations, extracted by a randomly initialized lightweight feature extractor $\psi_{\vartheta}$ (much smaller than the pretrained model to be unlearned). The resulting partition of the dataset can be written as:
$\begin{split}\{I_{p}\}_{p}\xrightarrow{\text{k-means partitioning}}\left\{\left\{I_{ij}\right\}_{j=1}^{n_{i}}\right\}_{i=1}^{k\times C}\end{split}$ (1)
@@
Algorithm 1 $k$-means Partitioning ($\mathbf{D}=\{I_{p}\}_{p}$, $\psi_{\vartheta}$)
@@
[…]
@@
4:  $\{I_{a}\}_{a=1}^{N_{c}}\leftarrow\{I_{p}\mid y_{p}=c\}$ $\triangleright$ Collect class images
5:  $\{\psi_{\vartheta}(I_{a})\}_{a=1}^{N_{c}}$ $\triangleright$ Compute feature vectors
6:  $\{\psi_{\vartheta}(I_{a})\}_{a=1}^{N_{c}}\xrightarrow{k\text{-means}}\bigl{\{}\{\psi_{\vartheta}(I_{b\kappa})\}_{b=1}^{n_{\kappa}}\bigr{\}}_{\kappa=1}^{k}$ $\triangleright$ Run $k$-means
7:  $\bigl{\{}\{\psi_{\vartheta}(I_{b\kappa})\}_{b=1}^{n_{\kappa}}\bigr{\}}_{\kappa=1}^{k}\to\bigl{\{}\{I_{b\kappa}\}_{b=1}^{n_{\kappa}}\bigr{\}}_{\kappa=1}^{k}$ $\triangleright$ Map to images
8:  $\mathcal{C}\leftarrow\mathcal{C}\cup\bigl{\{}\{I_{b\kappa}\}_{b=1}^{n_{\kappa}}\bigr{\}}_{\kappa=1}^{k}$ $\triangleright$ Add to global set
@@
[…]
@@
12:return $\{\{I_{ij}\}_{i=1}^{n_{j}}\}_{j=1}^{k\times C}$ $\triangleright$ Partitioned D dataset
@@
where $\psi_{\vartheta}$ is the feature extractor parameterized by $\vartheta$, and $\|\cdot\|$ denotes the Euclidean norm. The expectation is taken over a distribution $P_{\vartheta}$ of lightweight feature extractors to promote robustness, as adapted in [29].
@@ 3.3 Creating the Reduced Retain Dataset
Using Algorithm 2, we systematically separate the retain images (using partitioned training set and $\mathbf{F}$) into:
@@
$\mathbf{R}_{\text{reduced}}\;=\;\underbrace{\mathbf{Condense}\bigl{(}\mathbf{R}_{\text{free}}\bigr{)}}_{\text{condensed,\;low-importance}}\;\cup\;\underbrace{\mathbf{R}_{\text{residual}}}_{\text{uncondensed,\;high-importance}}.$
@@
[…]
@@ 3.4 Accelerated Approximate Unlearning
Carlini et al. [3] report that per-sample classification losses of deep networks are strongly non-Gaussian. They considered to “Gaussianise” losses by ad-hoc monotone maps (e.g. combination of $\exp$ and $\log$ maps), without guaranteeing that the output is normally distributed. We instead apply the probability–integral transform (PIT), i.e. for any continuous variable $X$ with CDF $F_{X}$, $Z=\Phi^{-1}\!\bigl{(}F_{X}(X)\bigr{)}$ is standard normal, where $\mathbf{\Phi^{-1}}$ is the probit (quantile) function of the standard normal distribution $\mathcal{N}(0,1)$ (mean 0, variance $1$). Substituting $F_{X}$ by a differentiable empirical surrogate yields an end-to-end differentiable pipeline whose transformed losses converge to $\mathcal{N}(0,1)$.
Following this motivation, we intend to transform arbitrary random samples, which may not be following Gaussian distribution towards samples following standard normal distribution. Let $x=(x_{1},\ldots,x_{N})\;\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\;F$, such that $F$ is an unknown continuous CDF. The empirical CDF at $x_{i}$ is $\hat{F}_{N}(x_{i})=\tfrac{1}{N}\sum_{j=1}^{N}\mathbf{1}\{x_{j}\leq x_{i}\}.$ Replacing the indicator by the temperature–controlled sigmoid $\sigma_{K}(z)=(1+e^{-Kz})^{-1},\;K>0,$ gives the differentiable estimator
@@
Because each entry of $\mathbf{Q}^{(K)}(x)$ lies in $(0,1)$, apply the probit function to obtain [22]
$\mathbf{z}^{(K)}=\mathbf{\Phi^{-1}}\!\bigl{(}\mathbf{Q}^{(K)}(x)\bigr{)}\in\mathbb{R}^{N}.$ (5)
Thus any sample $x$ from a continuous distribution is converted to a sample whose coordinates converge to i.i.d. standard normals as $K\to\infty$: $\mathbf{z}^{(K)}\;\xrightarrow{K\to\infty}\;\mathbf{\Phi^{-1}}\!\bigl{(}\hat{F}_{N}(x)\bigr{)}.$ A finite but large $K$ balances smooth gradients (for optimization) with true empirical CDF proximity.
@@ 3.4.1 Transforming losses of model to standard normal
Applying the vector transformation from (5) to the loss vectors yields $\mathbf{Z}_{F}^{(K)}=\mathbf{\Phi^{-1}}\!\bigl{(}\mathbf{Q}^{(K)}(\bm{\ell}_{F})\bigr{)},\qquad\mathbf{Z}_{T}^{(K)}=\mathbf{\Phi^{-1}}\!\bigl{(}\mathbf{Q}^{(K)}(\bm{\ell}_{T})\bigr{)},$ where $\mathbf{Q}^{(K)}$ is the differentiable CDF estimator from (4). As $K\to\infty$ the mapping converges to the empirical probit, and this allows use to play with transformed losses in distribution matching manner. As emphasize previously, practice we keep $K$ sufficiently large to preserve differentiability.
@@ 3.4.2 Membership-Inference Attack (MIA) Loss
[…]
@@ 3.4.3 Accelerated (A-) AMU Objective
Replacing per-sample $\mathcal{L}$ by its square $\mathcal{L}^{2}$ multiplies the per-sample gradient magnitude by $2\,\mathcal{L}$. Consequently, gradient descent updates make larger progress early in training, in such a way that large errors on $\mathbf{R}$ are penalized disproportionately, while small errors become less important in course of optimization. So the model departs more rapidly from the decision boundaries, which it learned earlier during training schedule. In order to prevent overfitting on high error retain dataset samples, we take square of average loss per batch, instead of per sample squaring. This also induces stability on choice of learning rate. Infact, in all of our experiments, we set a fixed learning rate for our approach on all models and datasets (including other hyperparameters), which is in strong contrast from other SOTA AMU methods that require hectic finetuning on each setting.
@@
Combining the controlled instability from steeper main loss, around decision boundaries, with reduction of training trajectories towards absolute unlearning solutions, forms a strong acceleration of unlearning. As would be seen later in our experiments, A-AMU objective minimization achieves competitive MIA suppression and retain accuracy after $\sim 1\,\text{epoch}$.
@@ 4.1 Experimental Setup
We evaluate our methods on CIFAR-10, SVHN, and CINIC-10 datasets using ResNet-18 and ResNet-50 models. Our approach is benchmarked against six AMU methods: Retraining, Catastrophic Forgetting (CF) (fine-tuning) [15, 13], Bad Distillation [6], $\mathbf{L_{1}}$-Sparsity [13], Pruning [13], and SCRUB [15]. For a fair comparison, implementations were adapted from their official repositories, and hyperparameters were rigorously fine-tuned for each model and dataset. [^2] Complete hyperparameter configurations are available in the ’hyperparameters’ folder of our code repository: https://github.com/algebraicdianuj/DC_Unlearning. We categorize these baselines into three tiers for analysis: naive (Retraining, CF), non-naive ($\mathbf{L_{1}}$-sparse, Pruning, SCRUB, Bad Distillation), and our proposed accelerated (A-) methods.
@@
Random Class Forgetting (Condensation Time for 1000 IPC: $47.7812\pm 1.3323$ s leading to 80.0% retain dataset reduction)
Retraining $48.61\pm 1.85$ $49.49\pm 2.38$ $100.00\pm 0.00$ $68.45\pm 2.51$ $\pm 0.00 0.00±0.00$$\pm 0.00$ $79.96\pm 1.10$ $56.46\pm 2.81$ $3580.71\pm 94.46$ $705.82\pm 0.30$
CF $51.07\pm 3.71$ $57.87\pm 2.74$ $100.00\pm 0.00$ $97.89\pm 0.18$ $0.00\pm 0.00$ $1.65\pm 1.73$ $81.28\pm 1.33$ $79.11\pm 1.89$ $1789.40\pm 47.15$ $352.38\pm 0.74$
Bad Distillation $57.97\pm 5.13$ $62.99\pm 6.81$ $7.31\pm 1.59 99$$9.69\pm 0.09$ $00\pm 0.00 0.0$$01\pm 0.03$ $77.68\pm 2.36$ $80.64\pm 1.16$ $1370.90\pm 34.32$ $572.67\pm 0.69$
$L_{1}$ Sparse $49.18\pm 2.20$ $58.95\pm 4.72$ $98.52\pm 0.55$ $98.54\pm 0.32$ $0.00\pm 0.00$ $4.03\pm 3.58$ $77.61\pm 2.12$ $79.61\pm 1.49$ $1446.43\pm 6.34$ $288.65\pm 0.31$
SCRUB $\pm 0.00 50.00±0.005$$\pm 0.00$ $5.50\pm 3.00 11$$1.14\pm 0.06$ $\pm 0.00 0.00±0.00$$\pm 0.00$ $13.88\pm 2.72$ $10.05\pm 0.09$ $759.90\pm 9.19$ $198.01\pm 0.21$
Pruning $48.37\pm 2.43$ $50.56\pm 0.81$ $7.10\pm 1.14 99$$9.08\pm 0.63$ $\pm 0.00 0.00±0.00$$\pm 0.00$ $76.85\pm 1.84$ $69.80\pm 3.61$ $722.06\pm 5.91$ $143.73\pm 0.23$
A-CF $52.10\pm 1.50$ $50.99\pm 0.67$ $97.75\pm 1.50$ $95.90\pm 2.73$ $0.00\pm 0.00$ $0.10\pm 0.19$ $77.03\pm 1.08$ $75.90\pm 3.82$ $218.30\pm 0.26$ $44.54\pm 1.80$
10% Uniform Forgetting (Condensation Time for 1000 IPC: $34.1780\pm 0.7403$ s leading to 39.8% retain dataset reduction)
@@
Unlearning performance is assessed across five key metrics: Test Accuracy ($\uparrow$), Retain Accuracy ($\uparrow$), Forget Accuracy ($\downarrow$), Unlearning Time ($\downarrow$), and a MIA Score ($\downarrow$) for privacy. The MIA score is derived from a powerful white-box LiRA attack [3], providing a robust measure of privacy risk. For our Blend method, the $k$-means partitioning feature extractor is a lightweight CNN, $35\times$ and $74\times$ smaller than ResNet-18 and ResNet-50 respectively, ensuring its preprocessing overhead remains minimal. All experiments were conducted on a single NVIDIA RTX 3070 GPU (8GB VRAM), except for experiments in Sections 4.4 and 4.5, which were performed on a laptop RTX 3070 (8GB VRAM) due to computational resource availability during the extended experimental phase.
@@
We categorize baselines into three tiers: naive (Retraining, CF), non-naive ($\mathbf{L_{1}}$-sparse, Pruning, SCRUB, Bad Distillation), and accelerated (A-), the last of which can contain acceleration of all naive and non-naive AMU methods.